Q. Solve for n: 3n-2/5=7/10

Answer by hit(1)

Your equation is 3n-2/5=7/10,
To solve this first we add 2/5 both side of the equation,then we will get,
3n - 2/5 + 2/5 = 7/10 + 2/5
3n = 7/10 + 2/5
3n = 7 /10 + 4/10
= (7 + 4)/10
= 11/10
= 1.1
Now we divide both side by 3,then we will get,
3n/3 = 1.1/3
n = 1.1/3
And that's all